OpenStudy (anonymous):

Solve the system of equations for one of the variables x+3y=6 x-3y=6

OpenStudy (anonymous):

a.x=5 b.y=6 c.x=6 d.x=3

OpenStudy (p0sitr0n):

6 =6 x+3y = x-3y 6y = x-x y=0 x +0 = 6 x =6

OpenStudy (anonymous):

Thanks=)